Solve the following system of linear inequalities by graphing.
x + 2y  3
2x – 3y  6
Both equations are less than or equal to.

Answer by checkley71(8403)

X+2Y<=3 MULTIPLY THIS EQUATION BY (-2) AND ADD TO THE SECOND EQUATION
2X-3Y<=6
-2X-4Y<=-6
-----------------------------------------
0-7Y<=0
Y<=0 SOLUTION
2X-3*0<=6
2X<=6
X<=6/2
X<=3 SOLUTION